OpenStudy (anonymous):

2x-y+z=-4 3x+y-2z=0 3x-y=-4 i have to solve it using elimination and i am so lost!

OpenStudy (anonymous):

2x-y+z=-4 (*2) 3x+y-2z=0 (the same) 3x-y=-4 (*-)

OpenStudy (anonymous):

So, 4x-2y+2z=-8 3x+y-2z=0 -3x+y=4

OpenStudy (anonymous):

Answer: x=-1 y=1 z=-1

OpenStudy (anonymous):

thanks so much